OpenStudy (anonymous):

Solve the system algebraically. 4x - 5y = 3 x = 3 - y What is the value of x?

OpenStudy (anonymous):

you can solve by substitution since x is by itself in one equation so plug (3 - y) in for x in the other equation 4(3 - y) - 5y = 3 12 - 4y - 5y = 3 12 - 9y = 3 -9y = -9 y = 1 plug 1 back in for y in one of the equations x = 3 - 1 x = 2 x = 2, y = 1 (2, 1) is the solution or intersection

OpenStudy (anonymous):

Correct me if i'm wrong :)